OpenStudy (anonymous):

Rewrite without rational exponents (-64xyz)^1/3

OpenStudy (anonymous):

Help??

OpenStudy (mathstudent55):

Can you write 64 as a number to a power?

OpenStudy (mathstudent55):

This takes care of the variables. \(\LARGE a^{\frac{1}{n}} = \sqrt[n]{a} \) For the number, first write 64 as a power. Then use this: \(\LARGE (a^m)^n = a^{mn} \)